# Detection of extended X-ray emission surrounding PSR B0656+14 with eROSITA

### v1

@@ Appendix A Point source contamination
PSR B0656+14 is slightly offset from the optical axis. To account for this, we adopt a position corresponding to an off-axis of $-5^{\prime}$ in both the $x$ and $y$ directions as a reference. The PSF images for this position from TM2, 3, 4 and 6 are combined to examine the fractions of flux contribution at various radii. At energies of 0.3, 1.5, and 3.0 keV, the flux contributions from the point source in the $3^{\prime}-4^{\prime}$ annulus are approximately 0.3%, 0.6%, and 0.8%, respectively. Beyond 4, the flux contribution become negligible according to the PSF model, with possible fluctuations $\lesssim 0.1$%. In the $4^{\prime}-6^{\prime}$ annulus, the flux of the non-thermal component is estimated to be $\sim$1.14of the central point source in 0.2-2.0 keV. Therefore, starting the innermost annulus at a radius of 4 is justified, ensuring that the fitting process is not contaminated by the central source.
@@
Note: Errors are at 90% confidence level. Columns from left to right are: annulus region, photon power-law index, flux in 0.2 - 2.0 keV, and fitting $\chi^{2}$ value divided by the number of degree-of-freedom.
@@
Appendix B Weak source correction
Fig. A4 show the surface brightness profiles of detected sources, for three flux bins. We convert the count rate to the flux of these point sources with the energy conversion factor (ECF) and the power-law model used by Brunner et al. (2022). One can see that for relatively bright sources, the detection efficiency is almost uniform in the field of view. However, for the weakest sources, the detection efficiency drops significantly beyond 16. The inefficiency of weak source detection may be the reason of the rising behavior at large radii as shown in Fig. A3. We adopt a correction method to correct such an effect. We fit the surface brightness profile from 4 to 16 using a constant, and the surface brightness profile beyond 16 using an exponential function. The difference between the constant surface brightness and the exponential function is taken as the correction values for fluxes measured beyond 16. The corrected flux distribution is shown in Fig. 1 of the main text, which gives nearly flat distribution at large radii.
Appendix C Multi-wavelength modelling
Using the pulsar halo model described in the main text, we calculate the ICS emission of the same population of electrons and positrons, and compare it with the HAWC data. The left panel of Fig. A5 shows the spectrum within $10^{\circ}$ radius of the pulsar, and the right panel shows the surface brightness profile between 8 and 40 TeV.

### v2

@@ Appendix A Point source contamination
PSR B0656+14 is slightly offset ($\sim 2.5^{\prime}$) from the optical axis. To provide a conservative upper bound on the expected PSF leakage, we adopt a position corresponding to an off-axis of $5^{\prime}$ in the $x$ and $0^{\prime}$ in the $y$ directions as a reference, where the $5^{\prime}$ is the minimum spatial interval for Shapelet PSF parameters. The Shapelet PSF images for this position from TM2, 3, 4 and 6 are combined to examine the fractions of flux contribution at various radii. At energies of 0.3, 1.5, and 3.0 keV, the flux contributions from the point source in the $3^{\prime}-4^{\prime}$ annulus are approximately 0.3%, 0.6%, and 0.7%, respectively. Beyond 4, the flux contribution become negligible according to the PSF model, with possible contribution $\lesssim 0.1$%. In the $4^{\prime}-6^{\prime}$ annulus, the flux of the non-thermal component is estimated to be $\sim$1.4% of the central point source in 0.2-2.0 keV. Therefore, starting the innermost annulus at a radius of 4 is justified, ensuring that the fitting process is not contaminated by the central source.
@@
Appendix B Validation of the PSF
The possible PSF leakage from the bright central source was one of the major concerns. To test the physical origin of the extended emission, we measured the 0.2–2.0 keV surface brightness profile of PSR B0656+14 after excluding all detectable point sources, correcting for vignetting, and normalizing to the innermost radius. We then compared its profile with multiple reference baselines, including both observations under the pointing mode and the survey mode models.
Figure A3: The eROSITA image of V* V702 CrA in 0.2–2 keV, corrected for vignetting and FWC. Two sectors show the sky regions used to extract the radial surface brightness profile, with a width of $2^{\prime}$. Small circles (in cyan and magenta) indicate the regions used to extract the spectra in Fig. A5.
The chosen reference source is the bright star V* V702 CrA (OBSID: 300017), which was observed in the same pointing mode. V* V702 CrA is a variable star of spectral type G2 located in the field of dark cloud TGUH2213P1 (Dobashi et al., 2005), a region with low background and sparse point sources. V* V702 CrA was observed at an off-axis angle of $\sim 4.5^{\prime}$, which is slightly bigger than the off-axis angle of PSR B0656+14 ($\sim 2.5^{\prime}$). Since the eROSITA PSF naturally degrades and broadens with increasing off-axis angle, the reference star’s PSF is expected to be intrinsically broader than that of the pulsar. Two reference regions of $90\arcdeg$ sectors (see Fig. A3), as well as the $360^{\circ}$ average, are chosen to derive the radial distributions around V* V702 CrA. The results are shown in Fig. A4. We find that the radial profiles for the two $90^{\circ}$ regions are slightly different, with the results of the green sector being more concentrated. The $360^{\circ}$ average profile lies in between of the two selected $90^{\circ}$ sectors. This may show that there is slight asymmetry of the PSF for the off-axis observations. However, the current data are limited in statistics, and the asymmetric effect cannot be clearly characterized. In any case, the observed radial profile of PSR B0656+14 is more extended than that of V* V702 CrA, even if the expected profile should be less extended if PSR B0656+14 is indeed a point source. This implies that extended emission around the pulsar exists. Quantitatively, the fraction of total 0.2–2.0 keV energy in the $4^{\prime}$$6^{\prime}$ annulus for V* V702 CrA is $(0.6\pm 0.2)\%$ as estimated using the east $90^{\circ}$ sector profile (in orange). For the other two profiles of V* V702 CrA, such a fraction should be even smaller. For PSR B0656+14, the corresponding fraction is $\sim 1.4\%$, significantly higher than the expectation from the V* V702 CrA profile.
For comparison, we also show in Fig. A4 the analytical 0.4–2.3 keV PSF model for the survey mode from Churazov et al. (2023) (red line) and the official source stacking PSF of 0.2–2.3 keV (grey line, retrieved from the eROSITA calibration results [^6] https://erosita.mpe.mpg.de/dr1/eROSITA_technical/flight_calibration and valid only within 4’; Merloni et al., 2024). We find that the pulsar’s profile is apparently similar to (slightly narrower than) these survey-mode PSF profiles. It may partly reflect a coincidence, given the current limitations in the PSF modelling that the source’s intrinsic extent broadens the profile to mimic the instrumental blurring in survey modes.
Figure A4: Comparison of the surface brightness profiles, normalized at $0.25\arcmin$. Blue: PSR B0656+14 (pointing mode); green and orange: V* V702 CrA (pointing mode; the 90° sectors with the same color as labelled in Fig. A3); purple: V* V702 CrA (pointing mode; 360° annulus); red: PSF model in the survey mode from Churazov et al. (2023); grey: source stacking PSF in survey mode from Merloni et al. (2024).
Spectra were also extracted from the portions of the $4^{\prime}$$8^{\prime}$ and $14^{\prime}$$18^{\prime}$ radial ranges within the east $90^{\circ}$ sector regions around V* V702 CrA, as shown in Fig. A5. The baseline spectral model consists of the cosmic X-ray background plus local bubble emission with absorption, i.e., $\tt tbabs*(powerlaw+apec)$, giving $\chi^{2}/\mathrm{dof}=369.02/322$. By comparing the two spectra, we find no significant difference in the foreground absorption between the two regions, suggesting that it does not affect the surface brightness profile. This result is consistent with the HI observation presented by HI4PI Collaboration et al. (2016). Adding an additional low-metallicity $\tt apec$ component with a fixed temperature $kT=0.83$ keV (corresponding to the emission from the central V* V702 CrA) does not improve the fit significantly (F-test: $F=3.466$, $p=0.0636$).
Figure A5: Simultaneous fitting of inner and outer regions around V* V702 CrA. The cyan and magenta lines correspond to the spectra extracted from the circular regions of the same colors in Fig. A3, respectively.
These results indicate that the leakage of emission from the central pulsar to the region beyond $4^{\prime}$ is very limited, and the excess around PSR B0656+14 is likely not due to the PSF contamination.
Appendix C Weak source correction
@@
The original surface brightness profile of the flux excesses is shown in Fig. A6, which gives a rising behavior at large radii. We expect that this issue should be due to the contamination of weak sources below the detector sensitivity. Fig. A7 show the surface brightness profiles of detected sources, for three flux bins. We convert the count rate to the flux of these point sources with the energy conversion factor (ECF) and the power-law model used by Brunner et al. (2022). One can see that for relatively bright sources, the detection efficiency is almost uniform in the field of view. However, for the weakest sources, the detection efficiency drops significantly beyond 16. We adopt a correction method to correct such an effect. We fit the surface brightness profile from 4 to 16 using a constant, and the surface brightness profile beyond 16 using an exponential function. The difference between the constant surface brightness and the exponential function is taken as the correction values for fluxes measured beyond 16. The corrected flux distribution is shown in Fig. 1 of the main text, which gives nearly flat distribution at large radii.
Appendix D Multi-wavelength modelling
Using the pulsar halo model described in the main text, we calculate the ICS emission of the same population of electrons and positrons, and compare it with the HAWC data. The left panel of Fig. A8 shows the spectrum within 10°radius of the pulsar, and the right panel shows the surface brightness profile between 8 and 40 TeV.